40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12028     Entry time: Thu Mar 10 03:03:11 2016
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Power stable, but no RF handoff 

[ericq, Gautam]

We worked on getting the DRFPMI back up and running, hoping the ALS performance was good enough. 

We did succeed in bringing in enough of the AO path to stabilize arm powers > 100, but failed at the full RF DARM handoff. 

REFL165 angle was adjusted to -86 to minimize PRCL in the Q signal. 

The AS110 signals are mysteriously huger than they used to be. Whitening gain reduced to 15dB from 27dB. Old trigger thresholds are still fine.

The new AUX X laser has a different sign for the temperature-> frequency coupling, so our usual convention of "beatnote goes up when temp slider goes up" meant the ALSX input matrix elements had to change sign.

We think the POPDC PD (which I think is the POP2F PD) may be miscentered, since in PRMI configuration, its maximum does not coincide with the REFLDC minimum, and leaves a sizeable TEM10 lobe on the REFL camera. This was a pain. 

ELOG V3.1.3-